# Conjuring Semantic Similarity

### v2

@@ Abstract
The semantic similarity between sample expressions measures the distance between their latent ‘meaning’. Such meanings are themselves typically represented by textual expressions, often insufficient to differentiate concepts at fine granularity. We propose a novel approach whereby the semantic similarity among textual expressions is based not on other expressions they can be rephrased as, but rather based on the imagery they evoke. While this is not possible with humans, generative models allow us to easily visualize and compare generated images, or their distribution, evoked by a textual prompt. Therefore, we characterize the semantic similarity between two textual expressions simply as the distance between image distributions they induce, or ‘conjure.’ We show that by choosing the Jensen-Shannon divergence between the reverse-time diffusion stochastic differential equations (SDEs) induced by each textual expression, this can be directly computed via Monte-Carlo sampling. Our method contributes a novel perspective on semantic similarity that not only aligns with human-annotated scores, but also opens up new avenues for the evaluation of text-conditioned generative models while offering better interpretability of their learnt representations.
@@ 1 Introduction
While measuring semantic similarity comes natural to humans who share significant knowledge and experience, defining semantic similarity for trained models is non-trivial since text can often be ambiguous or open to multiple (subjective) interpretations. On the other hand, visual elements often transcend communication barriers and allow for comparison at finer granularity. Comparing images is also relatively simple – unlike words, pixel values do not depend on distant knowledge or context.
Hence, instead of comparing images by the captions which describe them, we propose the converse: comparing textual expressions in terms of the images they conjure. In other words, we propose an expanded notion of meaning that is purely “visually-grounded”. This would be hard if not impossible for humans, since the process requires visualizing and comparing ‘mental images’ each individual can conceive, but it is straightforward for trained models.
We propose to leverage modern image generative models, in particular the class of text-conditioned diffusion models, for this purpose. By doing so, the semantic similarity between two text passages can simply be measured by the similarity of image distributions generated by a model conditioned on those passages.
There is a technical nugget that needs to be developed for the method to be viable, which is how to compare diffusions in the space of images. To address this, we propose to leverage the Jensen Shannon Divergence between the stochastic differential equations (SDEs) that govern the flow of the diffusion model, which we will show to be computable using a Monte-Carlo sampling approach.
We will show that our simple choice of distance already leads to results comparable to zero-shot approaches based on large language models. To validate our proposed definitions, we further run ablation studies over several components of our method, demonstrating robustness to specific choices of diffusion models and their inference algorithms.
To summarize the contributions of this work, we propose an approach for evaluating semantic similarity between text expressions that is grounded in the space of visual images. Our method has a unique advantage over traditional language-based methods that, in addition to providing a numerical score, it also provides a visual expression, or ‘explanation’, for comparison, enabling better interpretability of the learnt representations (Figure 1). Additionally, our method is the first to enable quantifying the alignment of semantic representations learnt by diffusion models compared to that of humans, which can open up new avenues for the evaluation text-conditioned diffusion models. Finally, we note that our approach can be used to derive many possible variants based on the metric used to compare images, which we leave for future exploration.
@@ 2 Related Works
The Distributional Hypothesis (Harris, 1954) forms the basis for statistical semantics, characterizing the meaning of linguistic items based on their usage distributions. This is also closely related to Wittgenstein’s use theory of meaning (Wittgenstein, 1953), often popularized as “meaning is use”. Many methods have been developed in machine learning and Natural Language Processing (NLP) literature to compute these semantic spaces, including Word2Vec (Mikolov et al., 2013). In light of modern Large Language Models (LLMs), Liu et al. (2023) defined the space of meanings for autoregressive models to be the distribution over model continuations for any input sequence. This has been used to define notions of semantic distances and semantic containment between textual inputs. Achille et al. (2024) further defined conceptual similarity between images by projecting them into the space of distributions over complexity-constrained captions, producing similarity scores that strongly correlate with human annotations. Soatto et al. (2023) generalizes these definitions by considering meanings as equivalence classes, where partitions can be induced by either an external agent or the model itself. Vector-based representations such as those obtained from CLIP Radford et al. (2021), or in general any sentence embedding model (Devlin et al., 2018; Opitz & Frank, 2022) have also been defined specifically for the computation of semantic distances. Such representations, however, are often difficult to interpret.
@@
Common metrics used to evaluate diffusion models are, among many others, the widely-used FID (Heusel et al., 2017) score, Kernel Inception Distance (Bińkowski et al., 2018), and the CLIP score (Hessel et al., 2021). Despite these choices, Stein et al. (2024) recently discovered that no existing metric used to evaluate diffusion models correlates strongly with human evaluations. While existing evaluations often focus on the quality and diversity of generations, our method is the first to evaluate semantic alignment of the representations learnt by diffusion models. Several techniques have also been developed to interpret the generation of diffusion models. Kwon et al. (2022); Park et al. (2023) edit the bottleneck representations within the U-Net architecture of diffusion models for semantic image manipulation. Gandikota et al. (2023) identifies low-rank directions corresponding to various semantic concepts, and Li et al. (2023) shows that diffusion models can be used for image classification. Kong et al. (2023a, b) frame diffusion models using information theory, improving interpretability of their learnt semantic relations. Orthogonal to these works, our method enables visualizing the semantic relations between textual prompts in natural language learnt by diffusion models via the distributions over their generated imageries.
@@ 3.2 Construction
return $d/k$triangleright Return similarity score
@@
[…]
Since our goal is to define a semantic distance between textual prompts $y_{1}$ and $y_{2}$ by comparing the distributions over images that they conjure, we can achieve this by computing a discrepancy function between the respective SDEs that they induce. In particular, we will use the Jensen–Shannon (JS) divergence, which is simply the symmetrized Kullback–Leibler (KL) divergence between two SDEs. In the following, we show how this divergence can be computed via a Monte-Carlo approach.
@@
[…]
Novikov’s Condition guarantees that this term is a Martingale so, following the approach of Song et al. (2021), the KL divergence between the two SDEs can be simplified as
[…]
@@
$d_{ours}(y_{1},y_{2}):=JSM(\mathbb{P}_{1}||\mathbb{P}_{2})=\frac{1}{2}D_{KL}(\mathbb{P}_{2}||\mathbb{P}_{1})+\frac{1}{2}D_{KL}(\mathbb{P}_{1}||\mathbb{P}_{2})$
@@ 4.2 Baselines
Direct output comparisons: For the same initial condition, we directly compute the difference between the images produced by the two different labels. This is defined as $d_{\text{output}}:=\mathbb{E}_{\bm{x}\sim\pi}\left[\|\psi_{\theta}(\bm{x}|y_{1})-\psi_{\theta}(\bm{x},|y_{2})\|_{2}^{2}\right]$, where $\psi_{\theta}(\cdot|y):\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ represents the full reverse diffusion process from the noise prior $\pi$ at time $T$ to the output distribution at time 00.
@@ 4.3 Empirical Validation
Table 1: Comparison with zero-shot methods on Semantic Textual Similarity benchmarks. ;$\ast$;$\dagger$;$\ddagger$ ; † ; ‡ indicate results taken from Ni et al. (2021); Gao et al. (2021); Liu et al. (2023) respectively. Expectedly, our zero-shot approach does not perform as well as embedding models such as CLIP (Radford et al., 2021) and SimCSE-BERT (Gao et al., 2021), which are trained specifically for semantic comparison tasks. Nevertheless, semantic structures extracted from text-conditioned diffusion models (StableDiffusion) using our method are still relatively aligned with that of human annotators, rivaling those extracted from autoregressive Large Language Models while outperforming encoder-based language models such as BERT (Devlin et al., 2018).
@@
To quantify this, we use the Semantic Textual Similarity (STS) (Agirre et al., 2012, 2013, 2014, 2015, 2016; Cer et al., 2017) and Sentences Involving Compositional Knowledge (SICK-R) (Marelli et al., 2014) datasets, containing pairs of sentences each labelled by human annotators with a semantic similarity score ranging from 0-5. We then use our method to compute the image-grounded similarity score, and measure their resulting Spearman Correlation with the annotations.
Interestingly, our experiments in Table 1 show that our visually-grounded similarity scores exhibit significant degrees of correlation with that annotated by humans. While, expectedly, our method presently lags behind embedding models trained specifically for semantic comparison tasks, we show that our visually-grounded similarity scores can rival that produced by existing large language models up to 33B in size. Since our work is the first to formalize and evaluate semantic alignment for this class of conditioned diffusion models, we have also included results from baselines with which our approach is directly comparable against. Our method convincingly outperforms all baseline methods for evaluating semantic similarity in diffusion models. As a remark, we note the large standard deviations in these scores across datasets for all methods, suggesting that what matters for validation purposes is that our distances correlate significantly with that of annotators, rather than obtaining “state-of-the-art” alignment scores since ‘ground truth’ is fundamentally subjective in each of these benchmarks.
@@
Figure 2: Qualitative evaluation of conjured semantic similarity. (Left) shows that nouns cluster based on shared hypernym classes: Dogs (puppy, poodle, dalmatian, pug) form a visible cluster in the top-left 4x4 block, while marine animals (whale, shark, dolphin, sealion) form another cluster in the bottom-right 4x4 block. (Right) shows that the same pattern holds for flying-related action verbs (elevate, ascend, soar, glide) v.s. negative stative verbs (disappoint, grieve, worry, regret).
@@ 4.4 Empirical Analysis
The computational feasibility of our method depends on the number of Monte-Carlo steps (i.e. $k$ in Algorithm 1) required to produce a reliable approximate of the desired distance. We ablate over choices of $k\in\{1,\ldots,5\}$ in Figure 3 and show that the deviation of scores when evaluated on the STS-B dataset is small (±$\pm 0.77$) across different choices of $k$. This finding is promising, as it implies that our method can be computationally efficient, requiring only a small number of Monte-Carlo iterations to converge.
@@ 5 Discussion and Limitations
Our method has several limitations. First, imageries might indeed not be sufficient to fully capture the meaning of certain expressions, such as mathematical abstractions (like ‘imaginary numbers’) and metaphysical concepts (like ‘conscience’). Furthermore, many modern diffusion models use a pre-trained text-encoder to pre-process textual prompts. This means that representation structures obtained from the diffusion model outputs would be bottle-necked by those learnt by the text encoder. However, this limitation can be mitigated by the development of better encoders, such as those based on LLMs (BehnamGhader et al., 2024). Additionally, these encoder-based (vector) representations are often difficult to interpret, while our proposed method offers a way to visualize and interpret the learnt semantic similarities between textual expressions. Lastly, computation costs also remain a key limitation of our method, since it requires several inference passes through the diffusion model to compute a single semantic similarity score, mitigated only partially by the conclusions of our ablation study on required number of iterations.
Nevertheless, our method is the first to show that textual representations can be meaningfully compared for diffusion models by “grounding” them in the space of conjured images. We introduce the first method for evaluating semantic alignment of text-conditioned diffusion models. In particular, our work enables not only qualifying (via visual ‘explanations’), but also quantifying the alignment of this resulting semantic space with that of human annotators. Our method also enable fine-grained analysis of the failure modes of existing diffusion models, pinpointing specific areas where they align poorly with human annotators. Our general framework of conjuring semantic similarity is applicable to the broader class of image generative models, and we leave the exploration of applications beyond diffusion models to future work.

### v4

@@ Abstract
The semantic similarity between sample expressions measures the distance between their latent ‘meaning’. These meanings are themselves typically represented by textual expressions. We propose a novel approach whereby the semantic similarity among textual expressions is based not on other expressions they can be rephrased as, but rather based on the imagery they evoke. While this is not possible with humans, generative models allow us to easily visualize and compare generated images, or their distribution, evoked by a textual prompt. Therefore, we characterize the semantic similarity between two textual expressions simply as the distance between image distributions they induce, or ‘conjure.’ We show that by choosing the Jeffreys divergence between the reverse-time diffusion stochastic differential equations (SDEs) induced by each textual expression, this can be directly computed via Monte-Carlo sampling. Our method contributes a novel perspective on semantic similarity that not only aligns with human-annotated scores, but also opens up new avenues for the evaluation of text-conditioned generative models while offering better interpretability of their learnt representations.
@@ 1 Introduction
While measuring semantic similarity comes natural to humans who share significant knowledge and experience, defining semantic similarity for trained models is non-trivial. For the class of Large Language Models (LLMs), Liu et al. (2023) defined the space of meanings for autoregressive models as the distribution over model continuations for any given input sequence. Semantic alignment of these models with humans can be directly quantified by comparing the geometric and algebraic properties of this space - for instance, measuring whether prompts semantically similar to humans are also closely embedded in this space of meanings.
In this work, we study the converse: Interpreting the learnt semantic space of image (rather than text) generation models, and how well they align with that of human annotators. Instead of comparing images by the captions that describe them, we compare textual expressions in terms of the images they conjure. In other words, we propose an expanded notion of meaning that is purely “visually-grounded”. This would be hard if not impossible for humans, since the process requires visualizing and comparing ‘mental images’ each individual can conceive, but it is straightforward for trained models.
We focus on the class of text-conditioned diffusion models. Under our definition, the semantic similarity between two text passages for a model can simply be measured by the similarity of image distributions generated by the model, conditioned on those passages.
There is a technical nugget that needs to be developed for the method to be viable, which is how to compare diffusions in the space of images. To address this, we propose to leverage the Jeffreys Divergence between the stochastic differential equations (SDEs) that govern the flow of the diffusion model, which we will show to be computable using a Monte-Carlo sampling approach.
Although our goal is primarily to introduce a method to quantify and visualize the semantic representations learnt by image generation models, we will show that our simple choice of distance already leads to results comparable to zero-shot approaches based on large language models. To validate our proposed definitions, we further conducted ablation studies on several components of our method, demonstrating robustness to the specific decoding algorithms used for multiple choices of diffusion models.
To summarize the contributions of this work, we propose an approach for evaluating semantic similarity between text expressions that is grounded in the space of visual images. Our method has a unique advantage over traditional language-based methods that, in addition to providing a numerical score, it also provides a visual expression, or ‘explanation’, for comparison, enabling better interpretability of the learnt representations (Figure 1). Additionally, our method is the first to enable quantifying the alignment of semantic representations learnt by diffusion models compared to that of humans, which can open up new avenues for the evaluation text-conditioned diffusion models.
@@ 2 Related Works
The Distributional Hypothesis (Harris, 1954) forms the basis for statistical semantics, characterizing the meaning of linguistic items based on their usage distributions. This is also closely related to Wittgenstein’s use theory of meaning (Wittgenstein, 1953), often popularized as “meaning is use”. Many methods have been developed in machine learning and Natural Language Processing (NLP) literature to compute these semantic spaces, including Word2Vec (Mikolov et al., 2013). In light of modern Large Language Models (LLMs), Liu et al. (2023) defined the space of meanings for autoregressive models to be the distribution over model continuations for any input sequence. This has been used to define notions of semantic distances and semantic containment between textual inputs. Achille et al. (2024) further defined conceptual similarity between images by projecting them into the space of distributions over complexity-constrained captions, producing similarity scores that strongly correlate with human annotations. Soatto et al. (2023) generalizes these definitions by considering meanings as equivalence classes, where partitions can be induced by either an external agent or the model itself. To evaluate generative distributions, Pillutla et al. (2023) quantifies the gap between target and model distributions by estimating the area under divergence frontiers in an embedding space. Vector-based representations such as those obtained from CLIP Radford et al. (2021), or in general any sentence embedding model (Devlin et al., 2018; Opitz & Frank, 2022) have also been defined specifically for the computation of semantic distances. Such representations, however, are often difficult to interpret.
@@
Common metrics used to evaluate diffusion models are, among many others, the widely-used FID (Heusel et al., 2017) score, Kernel Inception Distance (Bińkowski et al., 2018), and the CLIP score (Hessel et al., 2021). Despite these choices, Stein et al. (2024) recently discovered that no existing metric used to evaluate diffusion models correlates strongly with human evaluations. While existing evaluations often focus on the quality and diversity of generations, our method is the first to evaluate semantic alignment of the representations learnt by diffusion models. Several techniques have also been developed to interpret the generation of diffusion models. Kwon et al. (2022); Park et al. (2023) edit the bottleneck representations within the U-Net architecture of diffusion models for semantic image manipulation. Gandikota et al. (2023) identifies low-rank directions corresponding to various semantic concepts, and Li et al. (2023) shows that diffusion models can be used for image classification. Kong et al. (2023a; b) frame diffusion models using information theory, improving interpretability of their learnt semantic relations. Orthogonal to these works, our method enables visualizing the semantic relations between textual prompts in natural language learnt by diffusion models via the distributions over their generated imageries.
@@ 3.2 Construction
return $d/k$ $\triangleright$ Return similarity score
@@
We assume the standard conditions for existence and uniqueness of their solutions, in particular for all $t\in[0,T]$ we have $\mathbb{E}\left[\|\bm{x}(t)\|_{2}^{2}\right]\leq\infty$, and for all $y\in\mathcal{Y}$ and $x,x^{\prime}\in\mathbb{R}^{d}$, there exists constants $C,D$ such that $\|\mu_{\theta}(x,t,y)\|_{2}+\|g(t)\|_{2}\leq C(1+\|x\|_{2})$ and $\|\mu_{\theta}(x,t,y)-\mu_{\theta}(x^{\prime},t,y)\|_{2}\leq D\|x-x^{\prime}\|_{2}$. We further assume Novikov’s Condition holds for all pairs $y_{1},y_{2}\in\mathcal{Y}$, in particular we are guaranteed the following: $\mathbb{E}\left[\exp\left({\frac{1}{2}\int_{0}^{T}\|\mu_{\theta}(\bm{x},t,y_{2})-\mu_{\theta}(\bm{x},t,y_{1})\|_{2}^{2}dt}\right)\right]<\infty$.
Since our goal is to define a semantic distance between textual prompts $y_{1}$ and $y_{2}$ by comparing the distributions over images that they conjure, we can achieve this by computing a discrepancy function between the respective SDEs that they induce. In particular, we will use the Jeffreys divergence, which is simply the symmetrized Kullback–Leibler (KL) divergence between two SDEs. In the following, we show how this divergence can be computed via a Monte-Carlo approach.
@@
$\exp{\left(\int_{0}^{T}-\frac{1}{g(t)}\left(\mu_{\theta}(\bm{x},t,y_{1})-\mu_{\theta}(\bm{x},t,y_{2})\right)d\bar{\bm{w}}_{t}-\frac{1}{2}\int_{0}^{T}\frac{1}{g(t)^{2}}\|\mu_{\theta}(\bm{x},t,y_{1})-\mu_{\theta}(\bm{x},t,y_{2})\|_{2}^{2}dt\right)}$
Next, paralleling the derivation in Song et al. (2021), Novikov’s Condition guarantees that the stochastic integrand term $\frac{1}{g(t)}\left(\mu_{\theta}(\bm{x},t,y_{1})-\mu_{\theta}(\bm{x},t,y_{2})\right)$ is a Martingale. As a result, taking expectation under $\mathbb{P}_{2}$ would reduce it to zero, leaving only the drift term $\frac{1}{2}\int_{0}^{T}\frac{1}{g(t)^{2}}\|\mu_{\theta}(\bm{x},t,y_{1})-\mu_{\theta}(\bm{x},t,y_{2})\|_{2}^{2}dt$. Consequently, the KL divergence between the two SDEs can be simplified as
$D_{KL}(\mathbb{P}_{2}||\mathbb{P}_{1})$ $=\frac{1}{2}\mathbb{E}_{\mathbb{P}_{2}}\left[\int_{0}^{T}\frac{1}{g(t)^{2}}\|\mu_{\theta}(\bm{x},t,y_{1})-\mu_{\theta}(\bm{x},t,y_{2})\|_{2}^{2}dt\right]$
@@
$d_{ours}(y_{1},y_{2}):=D_{KL}(\mathbb{P}_{2}||\mathbb{P}_{1})+D_{KL}(\mathbb{P}_{1}||\mathbb{P}_{2})$
@@ 4.2 Baselines
Direct output comparisons: For the same initial condition, we directly compute the difference between the images produced by the two different labels. This is defined as $d_{\text{output}}:=\mathbb{E}_{\bm{x}\sim\pi}\left[\|\psi_{\theta}(\bm{x}|y_{1})-\psi_{\theta}(\bm{x},|y_{2})\|_{2}^{2}\right]$, where $\psi_{\theta}(\cdot|y):\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ represents the full reverse diffusion process from the noise prior $\pi$ at time $T$ to the output distribution at time 0.
@@ 4.3 Empirical Validation
Table 1: Comparison with zero-shot methods on Semantic Textual Similarity benchmarks, evaluated via Spearman Correlation (higher is better). $\ast;\dagger;\ddagger$ indicate results taken from Ni et al. (2021); Gao et al. (2021); Liu et al. (2023) respectively. Expectedly, our zero-shot approach does not perform as well as embedding models such as CLIP (Radford et al., 2021) and SimCSE-BERT (Gao et al., 2021), which are trained specifically for semantic comparison tasks. Nevertheless, semantic structures extracted from text-conditioned diffusion models (StableDiffusion) using our method are still well-aligned with human annotators, rivaling those extracted from autoregressive Large Language Models while outperforming encoder-based language models such as BERT (Devlin et al., 2018).
@@
To quantify this, we use the Semantic Textual Similarity (STS) (Agirre et al., 2012; 2013; 2014; 2015; 2016; Cer et al., 2017) and Sentences Involving Compositional Knowledge (SICK-R) (Marelli et al., 2014) datasets, containing pairs of sentences each labelled by human annotators with a semantic similarity score ranging from 0-5. We then use our method to compute the image-grounded similarity score, and measure their resulting Spearman Correlation with the annotations.
Interestingly, our experiments in Table 1 show that our visually-grounded similarity scores exhibit significant degrees of correlation with that annotated by humans. While, expectedly, our method presently lags behind embedding models trained specifically for semantic comparison tasks, we show that our visually-grounded similarity scores can rival that produced by existing large language models up to 33B in size. Since our work is the first to formalize and evaluate semantic alignment for this class of conditioned diffusion models, we have also included results from baselines with which our approach is directly comparable against. Our method convincingly outperforms all baseline methods for evaluating semantic similarity in diffusion models.
@@
Figure 2: Qualitative evaluation of conjured semantic similarity, visualized using pairwise distance matrices and hierarchial clustering. (Left) shows that nouns cluster based on shared hypernym classes: Dogs (puppy, poodle, dalmatian, pug) form a visible cluster in the bottom-right 4x4 block, while marine animals (whale, shark, dolphin, sealion) form another cluster in the top-left 4x4 block. (Right) shows that the same pattern holds for flying-related action verbs (elevate, ascend, soar, glide) v.s. negative stative verbs (disappoint, grieve, worry, regret).
@@ 4.4 Empirical Analysis
The computational feasibility of our method depends on the number of Monte-Carlo steps (i.e. $k$ in Algorithm 1) required to produce a reliable approximate of the desired distance. We ablate over choices of $k\in\{1,\ldots,5\}$ in Figure 3 and show that the deviation of scores when evaluated on the STS-B dataset is small ($\pm 0.77$) across different choices of $k$. This finding is promising, as it implies that our method can be computationally efficient, requiring only a small number of Monte-Carlo iterations to converge.
@@ 5 Discussion and Limitations
Our method has several limitations. Our work does not, nor is it meant to, provide a superior, general-purpose metric for computing semantic similarity (e.g. for applications such as retrieval). Indeed, imageries might indeed not be sufficient to fully capture the meaning of certain expressions, such as mathematical abstractions (like ‘imaginary numbers’) and metaphysical concepts (like ‘conscience’). Furthermore, many modern diffusion models use a pre-trained text-encoder to pre-process textual prompts. This means that representation structures obtained from the diffusion model outputs would be bottle-necked by those learnt by the text encoder. The meanings of textual expressions can often be ambiguous or open to multiple (subjective) interpretations. Our method does not resolve this linguistic ambiguity, since generated images are conditioned on the same potentially ambiguous text they are meant to clarify. Instead, it simply reveals which visual interpretation the model has associated with that expression.
Finally, computation costs also remain a key limitation of our method, since it requires several inference passes through the diffusion model to compute a single semantic similarity score, mitigated only partially by the conclusions of our ablation study on the required number of iterations.
Nevertheless, our method is the first to show that textual representations can be meaningfully compared for diffusion models by “grounding” them in the space of conjured images. We introduce the first method for evaluating semantic alignment of text-conditioned diffusion models. In particular, our work enables not only qualifying (via visual ‘explanations’), but also quantifying the alignment of this resulting semantic space with that of human annotators. Our method also enables fine-grained analysis of the failure modes of existing diffusion models, pinpointing specific areas where they align poorly with human annotators. Our general framework of conjuring semantic similarity is applicable to the broader class of image generative models, and we leave the exploration of applications beyond diffusion models to future work.
6 Acknowledgements
Research supported by ONR N00014-19-1-2229.
@@ Appendix A Additional Visualizations
Appendix B Additional Experiment Details
We detail the datasets used below:
[…]
Sentences Involving Compositional Knowledge (SICK-R) (Marelli et al., 2014): The SICK-R dataset contains 9.93K pairs of examples which involve compositional knowledge, and similarly provides a human-labeled semantic similarity score for each pair. Below contains an example:
Sentence A:
A man with a jersey is dunking the ball at a basketball game
Sentence B:
The ball is being dunked by a man with a jersey at a basketball game
Relatedness Score:
4.9
RG65(Rubenstein & Goodenough, 1965) contains human similarity ratings for 65 English noun pairs.
SimLex-999(Hill et al., 2015) contains 999 different concrete and abstract pairs of adjectives, nouns, and verbs. Each pair is also scored by annotators based on their semantic similarity.